50 ml of water sample requires 18 ml of 0.05M EDTA, during titration whereas 50 ml of




boiled water sample requires 12.5 ml of same EDTA in the titration. Calculate total,




permanent and temporary hardness of water

1
Expert's answer
2022-01-30T07:57:02-0500


50ml of water sample

=15ml of 0.01M EDTA

= ml of 0.01 EDTA

=300ml of 0.01M EDTA

=2×300ml of 0.01 N EDTA

=600 ml or 0.6 L of 0.01 eq. of CaCO3

=0.6×0.01×50g CaCO3 eq.

Then the total hardness=0.30g or 300mg of CaCO3 eq=300mg/L or ppm

now 50 ml of boiled water=5 ml of 0.01 M EDTA

1000ml of boiled water= ml of 0.01 M EDTA

=100 mL of 0.01 M EDTA

=200ml or 0.2 L of 0.01 N EDTA

=0.2×0.01×50 g of CaCO3 eq

=0.1 g or 100mg of CaCO3 eq

permanent hardness=100mg/L or ppm

permanent hardness=100mg/L or ppmTemporary hardness=300-100 = 200 ppm